### v1

# MD-Splatting: Learning Metric Deformation from 4D Gaussians in Highly Deformable Scenes

Abstract
Accurate 3D tracking in highly deformable scenes with occlusions and shadows can facilitate new applications in robotics, augmented reality, and generative AI. However, tracking under these conditions is extremely challenging due to the ambiguity that arises with large deformations, shadows, and occlusions. We introduce MD-Splatting, an approach for simultaneous 3D tracking and novel view synthesis, using video captures of a dynamic scene from various camera poses. MD-Splatting builds on recent advances in Gaussian splatting, a method that learns the properties of a large number of Gaussians for state-of-the-art and fast novel view synthesis. MD-Splatting learns a deformation function to project a set of Gaussians with non-metric, thus canonical, properties into metric space. The deformation function uses a neural-voxel encoding and a multilayer perceptron (MLP) to infer Gaussian position, rotation, and a shadow scalar. We enforce physics-inspired regularization terms based on local rigidity, conservation of momentum, and isometry, which leads to trajectories with smaller trajectory errors. MD-Splatting achieves high-quality 3D tracking on highly deformable scenes with shadows and occlusions. Compared to state-of-the-art, we improve 3D tracking by an average of 23.9 %, while simultaneously achieving high-quality novel view synthesis. With sufficient texture such as in scene 6, MD-Splatting achieves a median tracking error of 3.39 mm on a cloth of 1 $\times$ 1 meters in size. Project Website: https://md-splatting.github.io
Figure 1: We propose MD-Splatting, a method that improves state-of-the-art methods for simultaneous dynamic novel view synthesis and accurate 3D point tracking in highly deformable scenes. This figure shows the rendering and tracking of MD-Splatting in the six dynamic Blender [11] scenes used for evaluation. We will refer to the scenes in this Figure as Scenes 1, 2, 3, 4, 5 and 6 ordered from left to right.
1 Introduction
Three-dimensional tracking of clothing, towels, and similarly highly deformable objects as they drop, deform, and drape can facilitate new applications and improve the capabilities of existing applications. Generative AI and augmented reality applications can benefit from novel view synthesis for training, interaction, and scene editing [21]. Robots can benefit from tracking task-relevant geometries to dexterously manipulate deformable objects [46, 8, 47] and use novel-view synthesis to augment learning [20]. Simulation methods can improve fidelity using real2sim and related approaches that leverage real-world tracking data for parameter and method tuning [28]. However, tracking moving and morphing points poses many non-trivial challenges: points change color as they pass through shadows, occlusions interfere with the continuous visibility of points, and internal and external forces, such as material tension and air currents, cause points to move unpredictably.
Recent work demonstrated Gaussian splatting [25] can yield state-of-the-art novel view synthesis and rendering speeds exceeding 100 fps. 3D Gaussian Splatting fits the properties of a set of Gaussians, their color, position, and covariance matrix, using a fast differentiable rasterizer. An extension of 3D Gaussian splatting [33] showed dynamic scenes can be modeled by explicitly optimizing the properties of Gaussians over time, including the metric deformations. Metric deformations are the changes in the position of Gaussians in the world frame between two adjacent timesteps. Metric deformations across many timesteps yields 3D tracking of Gaussians, contrary to related work [42] with deformations available only in a canonical frame.
Explicitly optimizing the Gaussian pose as in Dynamic 3D Gaussians [33], the most relevant baseline to MD-Splatting, results in degraded performance with large deformations and shadows. Training time also scales linearly with the number of timesteps provided for each camera, meaning longer sequences take propotionally longer to train.
We propose MD-Splatting, a method that uses time-synchronized image frames from a calibrated multi-camera setup to track 3D geometries of deformable objects as they move through shadows and occlusions. MD-Splatting learns the canonical state of a set of Gaussians and a deformation function that maps the Gaussians into metric space. This enables tracking by recovering metric deformations, and novel-view rendering (through splatting) using a fast differentiable rasterizer.
We evaluate MD-Splatting in six photo-realistic synthetic scenes of varying difficulty. The scenes contain large deformations, shadows, and occlusions (see Figure 1 for visualization of the scenes and tracking trajectories by MD-Splatting). Empirical results show that MD-Splatting infers 16.7 % more accurate 3D tracking results compared to previous state-of-the-art [33], and achieves high-quality novel-view reconstruction results with an average PSNR of 39.1. In a scene with cameras at 1 m radius from the origin (i.e. Scene 6), MD-Splatting can track cloth deformation with as low as 3.39 mm median tracking error. To the best of our knowledge, MD-Splatting is the first approach for 3D tracking and view synthesis using canonical Gaussian splats.
In summary, we contribute:
  • The first approach that learns continuous metric deformations from a canonical space of Gaussians.
  • Experiments that suggest state-of-the-art performance in simultaneous 3D metric tracking and novel view synthesis. MD-Splatting improves tracking accuracy by an average of 16.7 % while achieving competitive view reconstruction.
  • A set of six synthetic scenes with large deformations, strong shadows, and occlusions. The scenes will be shared with the community.
2 Related Work
MD-Splatting builds on prior work in novel view synthesis in dynamic scenes to enable simultaneous 3D tracking and novel view synthesis. We discuss prior works in novel view synthesis in static and dynamic scenes, as well as methods for point tracking.
2.1 Neural Rendering for Novel View Synthesis
MD-Splatting builds on prior work in novel view synthesis, and uses photometric consistency as a signal to achieve 3D tracking. A popular novel view synthesis approach is NeRF [34], which uses neural networks to learn scene representations that are capable of photo-realistic novel view reconstruction. The learned networks render a 3D point by outputting its RGB radiance and density when viewed from an arbitrary camera view. NeRF then renders pixels using existing volume rendering techniques. Subsequent works improve NeRF along several axes: one is to speed up training and inference time via novel representations and system optimizations [36, 44, 31, 68, 50, 19, 32, 36, 35], or depth supervision [12, 1, 63, 37, 49]. Other works extend NeRF to more challenging conditions, such as sparser camera views [69, 10, 54, 39], fewer extrinsic camera calibrations [29, 66, 9, 24], transparent objects [26, 23] and reflective surfaces [57].
Particle-based methods use a more explicit representation than typical NeRF-based approaches. MD-Splatting builds on 3D Gaussian Splatting [25], which proposed a differential rasterizer to render a large number of Gaussian ‘splats,’ each with their state including color, position, and covariance matrix. Gaussian splatting achieves real-time rendering of novel views with state-of-the-art performance.
2.2 Dynamic Novel View Synthesis
The assumption of static scenes in neural rendering approaches prevents application to real-world scenarios with moving objects and people, such as the dynamic and deformable scenes in this work. One line of work to address this assumption is adding a time dimension to NeRF modeling [15, 18, 27, 65], either condition the neural field on explicit time input or a time embedding. Another line of work learns a deformation field to map into a canonical space [43, 41], i.e., every 4D point in space and time maps to a 3D point in a canonical NeRF. This approach can be improved via distinguishing between foreground and background [52], or via leveraging depth priors [1, 63, 37], and can achieve fast training [16, 53].
To accelerate the training of dynamic neural radiance fields, DeVRF [30] proposed to model the 3D canonical space and 4D deformation field of a dynamic, non-rigid scene with explicit and discrete voxel-based representations. To solve the overfitting issue resulting from a large number of grid parameters, DeVRF [30] introduces a static-to-dynamic learning paradigm to first reconstruct a 3D canonical voxel space using static multi-view images and then optimize a 4D deformation voxel space using few-view dynamic sequences. It achieves convergence of optimizing dynamic voxel radiance fields within $10$ minutes on a single NVIDIA GeForce RTX3090 GPU. However, since it relies on the static scene training stage to learn the scene information, it struggles to reconstruct dynamic scenes with complex deformations and shadows, especially with dynamic occlusions.
Several recent works extend the above approaches to 3D Gaussian splatting. Dynamic 3D Gaussians [33] explicitly model the position and covariance matrix of each Gaussian at each time step. This method struggles in dynamic scenes with large deformations, strong shadows, or occlusions. We build on another recent work, 4D Gaussian splatting [64], which uses feature encoding techniques proposed in HexPlanes [6] and K-planes [17], and learns a deformation field instead. To the best of our knowledge, all prior deformation-based work learned deformations in a non-metric canonical space without reconstructing metric deformations. We propose a method to learn metric deformations from a set of Gaussians with canonical properties, deformed to metric space using a neural-voxel deformation function. By training on adjacent timesteps, MD-Splatting allows enforcing physics-informed regularization terms to improve performance beyond photometric consistency.
2.3 Point Tracking
Point tracking methods, usually trained on large amounts of data, aided previous 3D tracking approaches by providing a strong prior [30]. We also construct several baselines that include point tracking methods (Section 6). Prior work on point tracking often studies tracking 2D points across video frames, where a dominant approach is training models on large-scale synthetic datasets containing ground-truth point trajectories [13, 14, 71, 22] or dense optical flows [60]. Optical flow [3, 51] or scene flow [56, 2, 58] can also be viewed as single-step point-tracking in 2D and 3D, respectively.
Another relevant line of work tightly couples dynamic scene reconstruction and motion estimation of non-rigid objects. A predominant setup is fusing RGBD frames from videos of dynamic scenes or objects [38]. Tracking or correspondence-matching methods see a progression from template-based tracking of objects with known shape or kinematics priors (such as human hand, face or body poses) [40, 45, 7], to more general shapes or scenes [72, 5, 4]. The main difference from these works is that we do not use depth input, and perform more rigorous quantitative evaluations on tracking specific points.
Most related to ours is the more recent methods that obtain tracking from neural scene rendering. DCT-NeRF[59] learns a coordinate-based neural scene representation that outputs continuous 3D trajectories across the whole input sequence. PREF [48] optimizes a dynamic space-time neural field with self-supervised motion prediction loss. Most recently, Luiten et al. [33] models dynamic 3D Gaussians explicitly across timestamps to achieve tracking. While our work also leverages 3D Gaussians, in contrary to the explicit modeling in Dynamic 3D Gaussians [33], we learn a deformation function that scales much better with video length, and we focus on deformable objects that are more challenging than the ball-throwing videos used in [33].
A main motivation for studying point tracking is the potential it can unlock for robotics applications: for example, RoboTAP [55] shows pre-trained point-tracking models improves sample efficiency of visual imitation learning.
3 Problem Statement
Given a set of timed image sequences captured from multiple cameras with known intrinsics and extrinsics, the objective is to learn a model that performs 3D tracking and novel view synthesis. Each image sequence is captured over the same time interval $t\in[0,H]$.
3D Tracking The 3D tracking objective is to recover the trajectory of any point in a dynamic scene by modeling metric deformations. Thus, the objective is to find a function $x_{t}=Q(x_{0},t_{0},t)$, where $x_{0}\in\mathbb{R}^{3}$ is the location of a point of interest at a chosen time $t_{0}$, while $x_{t}\in\mathbb{R}^{3}$ is the location of the same point at another chosen time $t$. The function $Q$ is valid for any point $x_{0}$ and any $t\in[0,H]$, allowing for tracking of any point in space.
Novel View Synthesis This objective is to recover novel views from arbitrary viewpoints. The extrinsics at any viewpoint can be captured by matrix $P$, with $P=K[R|T]$. Here $K$ is the intrinsics matrix, $R$ is the rotation matrix of a camera with respect to the world frame, and $T$ is the translation vector with respect to the world frame. The aim is to learn a function $V$ such that $I_{P,t}=V(P,t)$, where $I_{P,t}$ is an image rendered from a camera with extrinsics $P$ at time $t$. As with the tracking objective, the time parameter is valid for any $t\in[0,H]$.
4 Preliminary
4.1 Neural Radiance Fields (NeRF)
Neural radiance fields (NeRFs) trace rays through multiple camera views to train an MLP $\Phi_{r}$ to infer a radiance (RGB) and density ($\sigma)$ from a world coordinate $(x)$ and view direction $(\theta,\phi)$ inputs, which are typically concatenated with a positional encoding in the form of sinusoids with varying frequencies.
$\Phi_{r}(x,\theta,\phi)=(RGB\sigma)$ (1)
4.2 Gaussian Splatting
3D Gaussian Splatting [25] deploys a more explicit scene representation than NeRFs, by rendering a large set of Gaussians each defined by their mean position $\mu$ and covariance matrix $\Sigma$. Thus, for each $x\in\mathbb{R}^{3}$, its Gaussian $G(x)$ is
$G(x)=e^{-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)},$
Directly optimizing the covariance matrix $\Sigma$ would lead to infeasible covariance matrices, as they have to be positive semi-definite to have a physical meaning. Instead, Gaussian Splatting [25] proposes to decompose $\Sigma$ into a rotation $R$ and scale $S$ for each Gaussian:
$\Sigma=RSS^{T}R^{T},$
and optimize $R$, $S$, and the mean position.
Given the transformation W of a camera, the covariance matrix can be projected into image space as
$\Sigma^{\prime}=JW\Sigma W^{T}J^{T},$
where $J$ is the Jacobian of the affine approximation of the projective transformation.
During rendering, we compute the color $C$ of a pixel by blending $N$ ordered Gaussians overlapping the pixel :
$C=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$
where $c_{i}$ is the color of each point and $\alpha_{i}$ is given by evaluating a 2D Gaussian with covariance multiplied with a learned per-point opacity [67, 25].
4.3 Deformation Fields for Dynamic Scenes
Prior work showed that a deformation function combined with a static NeRF in a canonical space can enable novel view synthesis in dynamic scenes. The deformation function $F_{\mathrm{NeRF}}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ deforms a point in world coordinates $(x^{\prime})$ into a point in canonical space $(x)$. Prior work formulated $F_{\mathrm{NeRF}}$ as an MLP [42] and a multi-resolution voxel grid [16]. Wu et al. [64] applied a similar approach to arrive at Gaussian splatting of dynamic scenes. Given the state of a single canonical Gaussian, defined by $P=[\mu,S,R,\sigma,C]$ at time $t$, a deformation function takes the form
$P^{\prime}=F_{\mathrm{4DGS}}(P,t),$
where $F_{\mathrm{4DGS}}$, similar to the Hexplanes [6], contains a neural-voxel encoding in space and time. 4D-GS [64] starts with a coarse stage for initializing the canonical space, by setting $P^{\prime}$ = $P$, bypassing the deformation field and learning canonical properties directly. During the fine stage we learn the deformation function.
We propose MD-Splatting (Figure 2), based on 4D-Gaussians [64], to render novel views in dynamic scenes. The key differences with 4D-GS are: (1) we propose a simple method to track canonical Gaussians in 3D using a canonical $\rightarrow$ metric deformation function, (2) the output of the deformation function is different, and (3) using the method shown in Figure 3, we enforce regularization losses on the 3D trajectories of Gaussians.
Figure 2: MD-Splatting maps a set of Gausians with canonical properties to metric space using a deformation function $F$. The deformation function takes in the position of a Gaussian $(x,y,z)$ and a queried timestamp $t$, to infer shadow $s$, rotation $r^{\prime}$ and metric position $x^{\prime}$. During training, we use the metric positions and rotations to regularize the deformation function, considering the state at $t=\{i-1,i,i+1\}$ with Gaussian metric states $P^{\prime}_{t-1},P^{\prime}_{t},P^{\prime}_{t+1}$
5 Method
MD-Splatting achieves high-quality novel view synthesis using a canonical space of Gaussians and a deformation function to deform them to metric space (Section 5.1). To incentivize learning physically plausible deformations, MD-Splatting introduces several regularization terms (Section 5.2).
Figure 3: MD-Splatting uses three adjacent timesteps at every iteration to enforce physics-inspired regularization terms. All Gaussians are deformed to the metric space using the deformation function $F$, and rasterized to compute the photometric loss and its gradients. The positions of the Gaussians are used to compute the regularization terms based on local rigidity, isometry, and conservation of momentum (Section 5.2).
5.1 4D Gaussian Splatting
Canonical Neural Voxel Encoding. As with prior work, MD-Splatting learns a deformation function $F$ from a canonical space. We use a neural-voxel encoding to ensure $F$ has sufficient capacity to capture complex deformations. Prior work [17, 64, 16, 36] showed that neural-voxel encodings improve the speed and accuracy of novel view synthesis in dynamic scenes. We leverage HexPlanes [6, 64] to increase capacity for metric tracking and novel view synthesis simultaneously.
Figure 2 shows an overview of the canonical neural-voxel encoding. Each of the six voxel modules can be defined as $R(i,j)\in\mathbb{R}^{h\times lN_{i}\times lN_{j}}$. Here $\{{i,j}\}\in\{(x,y),(x,z),(y,z),(x,t),(y,t),(z,t)\}$, i.e., we adopt HexPlanes in all possible combinations. $h$ is the size of each feature vector in the voxel, $N_{i},N_{j}$ are the sizes of the HexPlanes in each dimension, $l$ is the upsampling scale. In every module, each plane has a different upsampling scale $l$. To query the multi-resolution voxel grids, we query each plane using bilinear interpolation to finally arrive at a feature vector used by the deformation MLP.
Deformation MLP. The deformation MLP takes in the voxel encoding and uses the encoding to deform the canonical Gaussians into metric space. Figure 2 shows the deformation MLP, which infers position, rotation, and a shadow scalar. We choose this set of outputs to model rigid-body transformations of each Gaussian and changes in illumination. Modeling changes in illumination is critical in the presence of shadows. The RGB color of each Gaussian is multiplied by the shadow scalar $s\in[0,1]$, and the shadow scalar is in the range $[0,1]$ by feeding the output of the MLP through a sigmoid activation function.
Next, we deform the Gaussians, modifying their mean positions $\mu$, rotation $R$, and color $C$, and arrive at a set of Gaussians in the metric space each with state $P$. The differentiable rasterizer from Gaussian Splatting [25] then renders the Gaussians to retrieve gradients for regressing both the canonical Gaussian states, and the parameters of the deformation function.
Unlike 4D-Gaussians [64], we propose to not infer opacity or scale using the deformation field. Optimizing for opacity and scale over time would allow Gaussians to disappear or appear instead of following the motion, which would make tracking less accurate. This design choice reduces the capacity of the deformation function, hence a lower view reconstruction quality as compared to 4D-Gaussians might be expected, see Table 2 for more information.
5.2 3D Tracking using 4D Gaussians
Motivation. Figure 3 shows the process of tracking Gaussians from the canonical space in metric space. By querying the deformation function $F$, we can track the position of a Gaussian along the entire trajectory.
Without additional supervision, this approach will not necessarily converge to metric deformations. Especially when objects include areas with little texture and uniform color, the solution space for all metric deformations is underconstrained by photometric consistency supervision. To better learn a metric deformation, we propose regularization terms inspired by physics.
After empirically evaluating several combinations of regularization terms, we adopt the rigidity and isometry losses proposed in Luiten et al. and add a momentum term. The first two terms capture a local rigidity loss and local isometry loss, which we compute based on the state of the $k$ nearest neighboring (KNN) Gaussians.
Local Rigidity Loss We assume the deformable material behaves linearly locally. With a large number of Gaussians, dense in 3D space, we expect nearest neighbors to move in similar directions and undergo similar rotations. Without this assumption, the trajectories of Gaussians are free to intersect and cross without a penalty. Instead, we introduce a local rigidity loss,
$\mathcal{L}_{i,j,t}^{\text{rigid}}=w_{i,j}\|(\mu_{j,t-1}-\mu_{i,t-1})-R_{i,t-1}R_{i,t}^{-1}(\mu_{j,t}-\mu_{i,t})\|_{2},$
where $\mu_{i,t}$ is the mean position of Gaussian $i$ at time $t$, $R_{i,t}$ is the rotation matrix of Gaussian $i$ at timestep $t$. This loss regularizes the change in position of the $k$ nearest neighbors as seen from Gaussian $i$, by transforming the current offset of the neighbors to the previous Gaussian orientation. We define
$w_{i,j}=\text{exp}\left(-\lambda_{w}\|\mu_{j,0}-\mu_{i,0}\|_{2}^{2}\right),$
an isotropic Gaussian used to decrease the weight of Gaussians at a greater distance from $\mu_{i,0}$. When a sufficiently large $\lambda_{w}$ is applied, and with a small enough $k$ for KNN, this loss fits the local linearity assumption. The total rigidity loss at time $t$ is
$\mathcal{L}^{\text{rigid}}_{t}=\frac{1}{k|\mathcal{P}|}\sum_{i\in\mathcal{P}}\sum_{j\in\text{knn}_{i}}\mathcal{L}_{i,j,t}^{\text{rigid}}.$
Here $\mathcal{P}$ is the set of all Gaussians.
Local Isometry Loss Another assumption we incentivize is to keep the relative position of the $k$ nearest neighbors constant w.r.t. $t=0$. With sufficient deformation, this assumption will be broken at a larger scale, but at a local scale, this regularization avoids drift from the ground truth trajectory. The isometry loss is
$\mathcal{L}_{t}^{\text{iso}}=\frac{1}{k|\mathcal{P}|}\!\sum_{i\in\mathcal{P}}\!\sum_{j\in\text{knn}_{i}}\!\!w_{i,j}\!\left|\left\|\mu_{j,0}{-}\mu_{i,0}\right\|_{2}-\left\|\mu_{j,t}{-}\mu_{i,t}\right\|_{2}\right|.$
Conservation of Momentum The final term we add is to incentivize the conservation of momentum. Newton’s first law states objects without external forces applied, given some mass $m$ and velocity vector $\mathbf{v}$, maintain their momentum $m\cdot\mathbf{v}$. We introduce the regularization term
$\mathcal{L}^{\text{momentum}}_{i,t}=\|\mu_{i,t+1}+\mu_{i,t-1}-2\mu_{i,t}\|_{1}.$
This term incentivizes a constant velocity vector and has the effect of imposing a low-pass filter on the 3D trajectories. It smooths out trajectories with many sudden changes of direction and magnitude (momentum). Table 1 shows the impact of the proposed regularization terms on tracking errors.
6 Experiments
We test the effectiveness of MD-Splatting on synthetic datasets of deformable objects. Section 6.1 provides details on our experiment setup, evaluation metrics, and baseline methods. Section 6.2 reports evaluation results from the compared methods and result analysis. Section 6.3 supplies additional ablation results on the training efficiency of MD-Splatting.
6.1 Experiment Setup
Metric Method Scene 1 Scene 2 Scene 3 Scene 4 Scene 5 Scene 6 Mean
3D MTE [mm] $\downarrow$ RAFT [51] ${}^{a}$ 89.687 151.363 295.66 135.468 125.554 67.861 144.266
RAFT Oracle [51]${}^{a}$${}^{b}$ 27.328 34.629 104.842 24.976 32.331 36.113 43.370
OmniMotion Oracle [60] ${}^{a}$${}^{b}$ 16.188 23.51 92.536 11.064 57.603 19.624 36.754
DeVRF [30] 147.293 78.487 449.74 123.745 77.736 133.391 168.399
DynaGS [33] 17.189 54.409 61.148 19.870 24.081 26.757 33.909
4D-GS [64] 7.788 93.409 118.692 16.422 7.033 4.600 41.324
MD-Splatting-NV 5.173 63.894 81.917 9.449 5.679 3.385 28.250
MD-Splatting-TK 2.881 46.257 88.169 9.686 4.741 3.175 25.818
3D $\delta_{\mathrm{avg}}~{}\uparrow$ RAFT [51] ${}^{a}$ 0.442 0.229 0.128 0.295 0.395 0.434 0.321
RAFT Oracle [51] ${}^{a}$${}^{b}$ 0.623 0.562 0.247 0.651 0.573 0.509 0.528
OmniMotion Oracle [60] ${}^{a}$${}^{b}$ 0.763 0.612 0.301 0.763 0.448 0.680 0.595
DeVRF [30] 0.344 0.368 0.094 0.326 0.471 0.398 0.334
DynaGS [33] 0.774 0.401 0.317 0.697 0.649 0.607 0.574
4D-GS [64] 0.787 0.337 0.257 0.700 0.817 0.883 0.630
MD-Splatting-NV 0.851 0.428 0.346 0.749 0.870 0.919 0.694
MD-Splatting-TK 0.912 0.511 0.338 0.780 0.903 0.939 0.731
3D Survival $\uparrow$ RAFT [51] ${}^{a}$ 0.935 0.891 0.738 0.919 0.807 0.891 0.864
RAFT Oracle [51] ${}^{a}$${}^{b}$ 1.000 1.000 1.000 1.000 0.999 0.974 0.996
OmniMotion Oracle [60] ${}^{a}$${}^{b}$ 1.000 0.993 1.000 0.991 0.870 1.000 0.976
DeVRF [30] 1.000 0.972 0.571 0.952 0.860 0.862 0.870
DynaGS [33] 1.000 1.000 0.958 1.000 1.000 1.000 0.993
4D-GS [64] 1.000 0.978 0.967 1.000 0.999 1.000 0.991
MD-Splatting-NV 0.999 1.000 0.987 1.000 1.000 1.000 0.998
MD-Splatting-TK 1.000 1.000 1.000 1.000 1.000 1.000 1.000
${}^{a}$ Method had access to ground truth depth. ${}^{b}$ Method had access to ground truth trajectories to pick the best camera view.
Table 1: 3D tracking results on the deformable cloth dataset (Figure 1). For each metric, the methods above the solid line had access to privileged information, see ${}^{a}$${}^{b}$ and Section 6.1 for more details. The results suggest that MD-Splatting outperforms the baselines in all averaged metrics. DynaGS [33] and OmniMotion [60] are both competitive but accumulate large tracking errors in scenes 1, 5, and 6. The results also show the regularization parameters improve the tracking performance compared to 4D-GS [64].
Dataset Preparation We use Blender to model dynamic cloth sequences and render photo-realistic images. We create 6 distinct scenes, each containing a different cloth with distinct visual and physical properties, and we render images from 100 different camera views and 40 consecutive time steps for training for a total of 4,000 images. The cloth deformations are introduced by dropping each cloth over one or a few invisible balls onto a ground plane or by constraining the cloth at an attachment point. We obtain ground-truth trajectories by tracking the mesh vertices of deformable objects in Blender.
Baselines. We compare to baseline methods: (1) 2D tracking approaches by projecting their inferences to 3D points using Blender’s ground truth depth and the camera’s ground truth projective geometry. While these methods were not designed for 3D tracking, they are well-known for their impressive 2D tracking performance. Their numbers aid in putting the tracking performance of the other baselines into context. We run RAFT [51] on all views and report the mean results as the RAFT model. We also provide two oracle methods which have access to privileged information: the ground truth trajectories. RAFT Oracle first evaluates on all views, to then output only the trajectories from the view with the lowest median trajectory error. We also report OmniMotion Oracle, which runs OmniMotion [60] on the frame with the lowest MTE for RAFT. Training OmniMotion takes roughly 12–13 hours on an Nvidia RTX 4090 GPU, making inference on all 100 views impractical. The numbers from RAFT Oracle and OmniMotion Oracle are not an apples-to-apples comparison with the other baselines, as to obtain their result they had to access privileged ground-truth trajectories.
(2) DeVRF [30] represents a dynamic, non-rigid scene with a 3D canonical voxel space and a 4D deformation voxel field. It proposes a 4D deformation cycle that models both the backward deformation from the deformed space to the canonical space and the forward deformation from the canonical space to the deformed space. Therefore, the point trajectories can be obtained by querying the forward deformation of canonical points at $t=0$ to every timestep’s deformed space through the 4D forward deformation space. While DeVRF was not designed for 3D tracking directly, it provides useful insight into how similar methods based on NeRF might perform in highly deformable scenes.
(3) Dynamic 3D Gaussians (DynaGS[33], which also builds upon 3D Gaussian splatting for dynamic novel view synthesis, except it explicitly models the positions and rotations of each Gaussian at each time-step. This results in straightforward tracking of any point via finding the trajectory of the learned Gaussian with the highest influence on the given point. Although the original paper assumes a known point cloud at the first frame, we instead use a randomly sampled point cloud for a fair comparison, with DynaGS and MD-Splatting both not using depth information. We also ran with a point cloud initialized using ground-truth depth from Blender but saw very little to no improvement.
(4) Finally, we compare to tracking using 4D-Gaussians [64] (4D-GS). We add our approach for tracking canonical Gaussian, as shown in Figure 3, to extract 3D trajectories from a learning view synthesis model. Comparing to 4D-GS serves to show the impact of the changes made in the model architecture and the regularization terms, to arrive at MD-Splatting.
Metrics Method Scene1 Scene2 Scene3 Scene 4 Scene5 Scene 6 Mean
[…]
DynaGS [33] 35.51 32.23 31.36 30.54 29.84 30.48 31.66
4D-GS [64] 40.76 40.14 44.38 42.05 33.70 34.83 39.31
MD-Splatting-NV 40.67 39.62 43.27 42.61 33.46 34.96 39.10
MD-Splatting-TK 38.76 38.16 40.34 40.61 32.47 32.39 37.12
[…]
DynaGS [33] 0.96 0.93 0.95 0.96 0.92 0.95 0.94
4D-GS [64] 0.99 0.97 0.98 0.99 0.96 0.98 0.98
MD-Splatting-NV 0.99 0.97 0.98 1.00 0.96 0.98 0.98
MD-Splatting-TK 0.99 0.97 0.98 0.99 0.95 0.97 0.98
[…]
DynaGS [33] 0.03 0.11 0.16 0.07 0.10 0.06 0.09
4D-GS [64] 0.01 0.07 0.07 0.02 0.07 0.02 0.04
MD-Splatting-NV 0.01 0.07 0.07 0.02 0.08 0.02 0.05
MD-Splatting-TK 0.02 0.08 0.08 0.03 0.08 0.04 0.05
Table 2: Novel view synthesis results on the deformable cloth dataset (Figure 1). The results suggest MD-Splatting significantly outperforms DynaGS [33] and DeVRF [30], while 4D-GS [64] performs slightly better. This is expected as a result of the additional regularization terms and deformation function design (Section 5). MD-Splatting only takes a minor decrease in view reconstruction performance, in order to significantly improve tracking. The results also suggest that MD-Splatting-TK performs worse in tracking as compared to MD-Splatting-NV, due to the stronger isometry regularization (Section 6.4).
[…]
We evaluate 3D tracking and novel-view synthesis results. For 3D tracking we evaluate each compared method on 1,000 randomly sampled points on each cloth. For novel-view synthesis, we evaluate using 200 unseen test-time cameras, and average the results across all cameras and all 40 time steps.
Figure 4: Qualitative novel view synthesis results from all baselines and MD-Splatting. Both 4D-GS [64] and MD-Splatting achieve the best qualitative results that are close to ground-truth renderings (especially on Scene 6, where the floral patterns on the cloth are reconstructed), albeit still fail to capture the fine-grained texture and wrinkles on Scene 2 and Scene3 cloth.
6.2 Main Results
3D Point Tracking. Following prior work [71, 33], we report median trajectory error (MTE), position accuracy ($\delta$), and the survival rate with a threshold of 0.5 [m] [33].
The results are summarized in Table 1. We make the following observations: (1) MD-Splatting outperforms the baselines in all averaged metrics, with a 23.9% better average MTE compared to the baseline methods. Dynamic 3D Gaussians [33], OmniMotion [60], and RAFT [51] are competitive but accumulate a significantly larger tracking error in scenes 1, 4, 5, and 6. For example, in scene 1, the baselines have a MTE of at least 170% higher as compared to MD-Splatting. Excluding our implementation of 4D-GS [64] for tracking, the clostest baseline is the OmniMotion [60] oracle model with a MTE 462% higher as compared to MD-Splatting. (2) MD-Splatting-TK performs better tracking compared to MD-Splatting NV, except for scene 3. The results suggest that setting $\lambda^{\text{iso}}=\text{1.0}$ leads to overall better tracking performance. Section 6.4 contains a more in-depth analysis on this parameter.
(3) MD-Splatting achieves the best performance in environments with textured cloth: scenes 1, 5, and 6. The texture makes photometric consistency a stronger supervision signal, leading to more accurate 3D tracking.
(4) The regularization parameters and change in deformation function architecture seem to improve the tracking performance compared to our implementation of 4D-GS [64]. MD-Splatting scores better metrics in all scenes. (5) The RAFT [51] and OmniMotion [60] results are highly dependent on the camera view angles, and we report the camera view with the best error values after evaluating all the test-time views; despite this given advantage, these methods still fail to achieve consistent and competitive tracking performance.
We hypothesize that part of the tracking error from Dynamic 3D Gaussians [33] is a result of converging Gaussian states to local optima. Especially with larger deformations and shadows, we imagine Gaussians could converge to local optima with relatively large tracking errors. The other source of error we hypothesize plays a role is the various regularization terms and their hyperparameters, which are predominantly focused on rigid motion.
(a) DeVRF [30]
(b) DynaGS [33]
(c) 4D-GS [64]
(d) MD-Splatting
Figure 5: Results on Scene 6: randomly sampled ground-truth trajectories in green, inferred trajectories in red, and the error in red. The differences between 4D-GS [64] are marked with green squares. Compared to the baseline methods, MD-Splatting results in fewer errors in 3D tracking.
Novel View Synthesis. We evaluate visual reconstruction quality. In Table 2, we report results on the three standard metrics commonly used in prior work: PSNR [61], SSIM [62], and LPIPS [70]. MD-Splatting-NV performs comparably to 4D Gaussians (4D-GS) on SSIM while obtaining marginally lower results on PSNR and LPIPS. DynaGS performs worse on view synthesis by a considerable margin. DeVRF obtains the lowest performance, and it does worse on the Scene 2 and 3 as compared to the other scenes—this suggests the textureless cloths impose additional challenges that the model cannot capture. MD-Splatting-TK yields a lower quality view reconstruction as compared to MD-Splatting-NV, as a result of the stronger isometry regularization (Section 6.4).
We jointly analyze results from 3D tracking 1 and novel view sythesis 2. We observe that 4D-GS under-performs DynaGS at 3D tracking in median trajectory error, but outperforms DynaGS at novel view synthesis. MD-Splatting is designed to achieve the best of both worlds. While close to 4D-GS in nature, MD-Splatting can track deformations better than DynaGS while keeping the high-quality novel-view synthesis capability.
Qualitative Results: Tracked Trajectories Figure 5 shows a quantitative comparison for tracking performance in Scene 6, overlaid on each method’ renders. The randomly sampled trajectories show examples of trajectories where DeVRF [30] and Dynamic 3D Gaussians [33] diverge from the ground truth. For DeVRF [30], we hypothesize a lack of high-quality view reconstruction is the cause of poor tracking, whereas Dyamic 3D Gaussians [33] has high-quality view reconstruction but has a few trajectories diverge from the ground truth. The difference between 4D-GS and MD-Splatting-NV is more subtle, and can be observed in the areas marked with green boxes. The green boxes highlight areas where 4D-GS [64] struggles to maintain tracking, whereas MD-Splatting-NV achieves better tracking.
Qualitative Results: Novel View Synthesis. Figure 4 shows a qualitative comparison for scenes 2, 3, and 6. Both 4D-GS and MD-Splatting-NV achieve the best qualitative results that are close to ground-truth (notably on Scene 6, where the floral patterns on the cloth are reconstructed), albeit still fail to capture the fine-grained texture and wrinkles on Scene 2 and 3 cloth. DynaGS also achieves high-quality rendering overall but display a blurry visual effect around the deformed (i.e., wrinkled) areas. We hypothesize that regressing Gaussian positions and orientations is more unstable, and prone to converge to local optima. DeVRF nearly completely fails in Scene 3 due to the lack of textures: the model overfits to one small patch on the cloth and ignores the rest of the scene; it also shows difficulty in learning the large deformations on edges, i.e. only capturing colors around the center of the cloth.
6.3 Training Efficiency Ablation
We compare the training efficiency between MD-Splatting-NV and Dynamic 3D Gaussians [33] when decreasing the number of time steps provided from each camera view. We evaluate on Scene 5 and report PSNR on the subsampled frames, training time, and median trajectory error in Table 3. We observe that for MD-Splatting, training time is less affected by the number of frames from each camera view, as compared to Dynamic 3D Gaussians [33]. This is a result of its approach to train one time step at a time. The results also suggest that MD-Splatting has more robustness to larger deformations between frames, leading to a better PSNR and median trajectory error.
Metric Method 40 Steps 20 Steps 14 Steps 10 Steps
[…]
MD-Splatting 33.47 33.00 32.80 32.91
Training DynaGS [33] 94.96 81.71 71.43 61.30
Time [mins] $\downarrow$ MD-Splatting 35.68 21.12 20.21 23.76
MTE [mm]$\downarrow$ DynaGS [33] 28.48 43.04 65.74 110.35
MD-Splatting 5.789 6.725 8.876 7.876
Table 3: Ablation on the number of time steps provided for each camera view. Fewer time steps means a lower time resolution.
[…]
6.4 Isometry Loss Ablation
[…]
7 Conclusions
In this work, we address the challenging problem of simultaneous 3D point-tracking and novel view synthesis in dynamic scenes with highly deformable objects. We introduced MD-Splatting, the first approach that learns continuous metric deformations from a canonical space of Gaussians. We empirically demonstrate that MD-Splatting outperforms baseline methods and achieves both high-quality dynamic scene reconstruction and high-accuracy 3D tracking on highly deformed cloth objects with occlusions and shadows. We also contribute a dataset of six synthetic scenes to facilitate future research.
Limitations and future work. MD-Splatting, similar to prior work on dynamic novel view reconstruction, requires a setup of multiple synchronized and calibrated cameras, which may require a significant engineering effort in real-world scenarios. This work also focused on scenes with only a cloth as the visible object, more scenes, simulation or real, with soft objects in a larger environment would be helpful. These limitations point to promising directions for future research.

### v2

@@
-MD-Splatting: Learning Metric Deformation from 4D Gaussians in Highly Deformable Scenes
+DeformGS: Scene Flow in Highly Deformable Scenes for Deformable Object Manipulation
 Abstract
-Accurate 3D tracking in highly deformable scenes with occlusions and shadows can facilitate new applications in robotics, augmented reality, and generative AI.
-However, tracking under these conditions is extremely challenging due to the ambiguity that arises with large deformations, shadows, and occlusions.
-We introduce MD-Splatting, an approach for simultaneous 3D tracking and novel view synthesis, using video captures of a dynamic scene from various camera poses.
-MD-Splatting builds on recent advances in Gaussian splatting, a method that learns the properties of a large number of Gaussians for state-of-the-art and fast novel view synthesis.
-MD-Splatting learns a deformation function to project a set of Gaussians with non-metric, thus canonical, properties into metric space.
+Teaching robots to fold, drape, or reposition deformable objects such as cloth will unlock a variety of automation applications.
+While remarkable progress has been made for rigid object manipulation, manipulating deformable objects poses unique challenges, including frequent occlusions, infinite-dimensional state spaces and complex dynamics.
+Just as object pose estimation and tracking have aided robots for rigid manipulation, dense 3D tracking (scene flow) of highly deformable objects will enable new applications in robotics while aiding existing approaches, such as imitation learning or creating digital twins with real2sim transfer.
+We propose DeformGS, an approach to recover scene flow in highly deformable scenes, using simultaneous video captures of a dynamic scene from multiple cameras.
+DeformGS builds on recent advances in Gaussian splatting, a method that learns the properties of a large number of Gaussians for state-of-the-art and fast novel-view synthesis.
+DeformGS learns a deformation function to project a set of Gaussians with canonical properties into world space.
 The deformation function uses a neural-voxel encoding and a multilayer perceptron (MLP) to infer Gaussian position, rotation, and a shadow scalar.
-We enforce physics-inspired regularization terms based on local rigidity, conservation of momentum, and isometry, which leads to trajectories with smaller trajectory errors.
-MD-Splatting achieves high-quality 3D tracking on highly deformable scenes with shadows and occlusions.
-Compared to state-of-the-art, we improve 3D tracking by an average of 23.9 %, while simultaneously achieving high-quality novel view synthesis.
-With sufficient texture such as in scene 6, MD-Splatting achieves a median tracking error of 3.39 mm on a cloth of 1 $\times$ 1 meters in size.
-Project Website: https://md-splatting.github.io
-Figure 1: We propose MD-Splatting, a method that improves state-of-the-art methods for simultaneous dynamic novel view synthesis and accurate 3D point tracking in highly deformable scenes.
-This figure shows the rendering and tracking of MD-Splatting in the six dynamic Blender [11] scenes used for evaluation.
+We enforce physics-inspired regularization terms based on conservation of momentum and isometry, which leads to trajectories with smaller trajectory errors.
+We also leverage existing foundation models SAM and XMEM to produce noisy masks, and learn a per-Gaussian mask for better physics-inspired regularization.
+DeformGS achieves high-quality 3D tracking on highly deformable scenes with shadows and occlusions.
+In experiments, DeformGS improves 3D tracking by an average of 55.8 % compared to the state-of-the-art.
+With sufficient texture, DeformGS achieves a median tracking error of 3.3 mm on a cloth of 1.5 $\times$ 1.5 m in area.
+Website: https://deformgs.github.io
+Keywords:
+Perception, Machine Learning in Robotics , Manipulation & Grasping
+1 Introduction
+Figure 1: We propose DeformGS, a method that improves state-of-the-art methods for accurate 3D point tracking in highly deformable scenes.
+This figure shows the rendering and tracking of DeformGS in the six dynamic Blender [9] scenes used for evaluation.
 We will refer to the scenes in this Figure as Scenes 1, 2, 3, 4, 5 and 6 ordered from left to right.
-1 Introduction
-Three-dimensional tracking of clothing, towels, and similarly highly deformable objects as they drop, deform, and drape can facilitate new applications and improve the capabilities of existing applications.
-Generative AI and augmented reality applications can benefit from novel view synthesis for training, interaction, and scene editing [21].
-Robots can benefit from tracking task-relevant geometries to dexterously manipulate deformable objects [46, 8, 47] and use novel-view synthesis to augment learning [20].
-Simulation methods can improve fidelity using real2sim and related approaches that leverage real-world tracking data for parameter and method tuning [28].
-However, tracking moving and morphing points poses many non-trivial challenges: points change color as they pass through shadows, occlusions interfere with the continuous visibility of points, and internal and external forces, such as material tension and air currents, cause points to move unpredictably.
-Recent work demonstrated Gaussian splatting [25] can yield state-of-the-art novel view synthesis and rendering speeds exceeding 100 fps.
-3D Gaussian Splatting fits the properties of a set of Gaussians, their color, position, and covariance matrix, using a fast differentiable rasterizer.
-An extension of 3D Gaussian splatting [33] showed dynamic scenes can be modeled by explicitly optimizing the properties of Gaussians over time, including the metric deformations.
-Metric deformations are the changes in the position of Gaussians in the world frame between two adjacent timesteps.
-Metric deformations across many timesteps yields 3D tracking of Gaussians, contrary to related work [42] with deformations available only in a canonical frame.
-Explicitly optimizing the Gaussian pose as in Dynamic 3D Gaussians [33], the most relevant baseline to MD-Splatting, results in degraded performance with large deformations and shadows.
-Training time also scales linearly with the number of timesteps provided for each camera, meaning longer sequences take propotionally longer to train.
-We propose MD-Splatting, a method that uses time-synchronized image frames from a calibrated multi-camera setup to track 3D geometries of deformable objects as they move through shadows and occlusions.
-MD-Splatting learns the canonical state of a set of Gaussians and a deformation function that maps the Gaussians into metric space.
-This enables tracking by recovering metric deformations, and novel-view rendering (through splatting) using a fast differentiable rasterizer.
-We evaluate MD-Splatting in six photo-realistic synthetic scenes of varying difficulty.
-The scenes contain large deformations, shadows, and occlusions (see Figure 1 for visualization of the scenes and tracking trajectories by MD-Splatting).
-Empirical results show that MD-Splatting infers 16.7 % more accurate 3D tracking results compared to previous state-of-the-art [33], and achieves high-quality novel-view reconstruction results with an average PSNR of 39.1.
-In a scene with cameras at 1 m radius from the origin (i.e.
-Scene 6), MD-Splatting can track cloth deformation with as low as 3.39 mm median tracking error.
-To the best of our knowledge, MD-Splatting is the first approach for 3D tracking and view synthesis using canonical Gaussian splats.
-In summary, we contribute:
-  • The first approach that learns continuous metric deformations from a canonical space of Gaussians.
-  • Experiments that suggest state-of-the-art performance in simultaneous 3D metric tracking and novel view synthesis.
-MD-Splatting improves tracking accuracy by an average of 16.7 % while achieving competitive view reconstruction.
-  • A set of six synthetic scenes with large deformations, strong shadows, and occlusions. The scenes will be shared with the community.
+Recent advances in robot learning have demonstrated impressive performance on challenging tasks, including rigid and deformable object manipulation.
+Scaling these approaches to deployment will require an improvement in robustness and learning from few demonstrations.
+A promising avenue for improving in robot learning performance are intermediate representations and foundation models, including 6D object pose estimation [10, 11, 27, 42, 50, 51, 55], semantic latent features [35], and 2D pixel-wise tracking [21, 52].
+However, perception and representations that will lead to robust manipulation of deformable objects remains an open challenge, due to self-occlusions, shadows, and varying (or lack of) textures.
+Three-dimensional dense point tracking, or 3D scene flow, is a useful representation for robot manipulation, as it provides flexibility to represent high-dimensional dynamic state changes, while the deformable objects drops, deforms, and drapes during manipulation.
+In particular, dense 3D scene flow can be an input to imitation learning policies [52, 3], can be used to learn a transition model [40], to identify and track task-relevant key points, or to create a digital twin through real2sim transfer.
+Recent work in monocular tracking has seen improvements in performance on datasets such as TAP-Vid [13], but it remains unclear how to effectively lift from 2D tracking to 3D for robotic spatial understanding in challenging highly deformable scenes.
+To overcome these limitations, Gaussian Splatting provides a promising avenue.
+Recent work demonstrated Gaussian Splatting [22, 23] can yield state-of-the-art novel-view synthesis and rendering speeds exceeding 100 fps.
+Concretely, 3D Gaussian Splatting uses a fast differentiable renderer to fit the colors, positions, and covariances of a set of Gaussians.
+An extension of 3D Gaussian splatting [29] showed dynamic scenes can be modeled by explicitly optimizing the properties of Gaussians over time, resulting in novel-view synthesis and scene flow.
+Explicitly optimizing the Gaussian pose as in Dynamic 3D Gaussians [29] may result in degraded performance with large deformations and shadows.
+The Gaussian properties may converge to local optima, especially in scenes with large deformations, strong shadows, and occlusions.
+We propose DeformGS, a method that uses time-synchronized image frames from a calibrated multi-camera setup to track 3D geometries of deformable objects as they move through shadows and occlusions.
+DeformGS learns the canonical state of a set of Gaussians and a deformation function that maps the Gaussians into world space.
+This enables tracking by recovering scene flow, and novel-view rendering (through splatting) using a fast differentiable rasterizer.
+We evaluate DeformGS in six photo-realistic synthetic scenes of varying difficulty.
+The scenes contain large deformations, shadows, and occlusions (Figure 1 shows the scenes and tracking trajectories computed by DeformGS).
+Empirical results show that DeformGS infers 55.8 % more accurate 3D tracking results compared to previous state-of-the-art [53, 29].
+In a scene with a 1.5 m $\times$ 1.5 m cloth (i.e., Scene 1), DeformGS can track cloth deformation with as low as 3.3 mm median tracking error.
+We also evaluate DeformGS in the real world on the Robo360 [26] dataset.
+We show qualitative results for tracking rigid and deformable objects in cluttered scenes, and study two robotics applications: (1) real2sim transfer to create a digital twin, and (2) tracking task-relevant keypoints for downstream grasping applications.
+In summary, our contributions are as follows:
+  • We provide the first approach designed to accurately perform 3D dense tracking for deformable objects using 4D Gaussians.
+  • We provide experiments that suggest state-of-the-art performance in simultaneous 3D metric tracking and novel view synthesis.
+DeformGS improves tracking accuracy by an average of 55.8% in synthetic experiments and demonstrates robust 3D tracking in the real world for deformable objects.
+The latter can be exploited as a representation for imitation learning and represents a new method for building digital twins.
+  • A set of six synthetic scenes with large deformations, strong shadows, and occlusions. We will open-source the scenes and as well as the source code.
 2 Related Work
-MD-Splatting builds on prior work in novel view synthesis in dynamic scenes to enable simultaneous 3D tracking and novel view synthesis.
-We discuss prior works in novel view synthesis in static and dynamic scenes, as well as methods for point tracking.
 2.1 Neural Rendering for Novel View Synthesis
-MD-Splatting builds on prior work in novel view synthesis, and uses photometric consistency as a signal to achieve 3D tracking.
-A popular novel view synthesis approach is NeRF [34], which uses neural networks to learn scene representations that are capable of photo-realistic novel view reconstruction.
-The learned networks render a 3D point by outputting its RGB radiance and density when viewed from an arbitrary camera view.
-NeRF then renders pixels using existing volume rendering techniques.
-Subsequent works improve NeRF along several axes: one is to speed up training and inference time via novel representations and system optimizations [36, 44, 31, 68, 50, 19, 32, 36, 35], or depth supervision [12, 1, 63, 37, 49].
-Other works extend NeRF to more challenging conditions, such as sparser camera views [69, 10, 54, 39], fewer extrinsic camera calibrations [29, 66, 9, 24], transparent objects [26, 23] and reflective surfaces [57].
+DeformGS builds on prior work in novel-view synthesis, and uses photometric consistency as a signal to achieve 3D tracking.
+A popular novel view synthesis approach is NeRF [30].
+It uses neural networks to learn scene representations that are capable of photo-realistic novel view reconstruction.
 Particle-based methods use a more explicit representation than typical NeRF-based approaches.
-MD-Splatting builds on 3D Gaussian Splatting [25], which proposed a differential rasterizer to render a large number of Gaussian ‘splats,’ each with their state including color, position, and covariance matrix.
-Gaussian splatting achieves real-time rendering of novel views with state-of-the-art performance.
+DeformGS builds on 3D Gaussian Splatting [22, 23] which belongs to the latter caregory.
+[22] proposed a differential rasterizer to render a large number of Gaussian ‘splats,’ each with their state including color, position, and covariance matrix.
+Contrary to the NeRF-based approaches, Gaussian splatting achieves real-time rendering of novel views with state-of-the-art performance.
 2.2 Dynamic Novel View Synthesis
-The assumption of static scenes in neural rendering approaches prevents application to real-world scenarios with moving objects and people, such as the dynamic and deformable scenes in this work.
-One line of work to address this assumption is adding a time dimension to NeRF modeling [15, 18, 27, 65], either condition the neural field on explicit time input or a time embedding.
-Another line of work learns a deformation field to map into a canonical space [43, 41], i.e., every 4D point in space and time maps to a 3D point in a canonical NeRF.
-This approach can be improved via distinguishing between foreground and background [52], or via leveraging depth priors [1, 63, 37], and can achieve fast training [16, 53].
-To accelerate the training of dynamic neural radiance fields, DeVRF [30] proposed to model the 3D canonical space and 4D deformation field of a dynamic, non-rigid scene with explicit and discrete voxel-based representations.
-To solve the overfitting issue resulting from a large number of grid parameters, DeVRF [30] introduces a static-to-dynamic learning paradigm to first reconstruct a 3D canonical voxel space using static multi-view images and then optimize a 4D deformation voxel space using few-view dynamic sequences.
-It achieves convergence of optimizing dynamic voxel radiance fields within $10$ minutes on a single NVIDIA GeForce RTX3090 GPU.
-However, since it relies on the static scene training stage to learn the scene information, it struggles to reconstruct dynamic scenes with complex deformations and shadows, especially with dynamic occlusions.
+The assumption of static scenes in neural rendering approaches prevents application to real-world scenarios with moving objects or humans, such as the dynamic and deformable scenes in this work.
+One line of work to address this assumption is adding a time dimension to NeRF modeling [15, 18, 25, 54].
+Prior works either condition the neural field on explicit time input or a time embedding.
+Another line of work learns a deformation field to map 4D points into a canonical space [37, 36], i.e., every 4D point in space and time maps to a 3D point in a canonical NeRF.
+DeVRF [28] proposed to model the 3D canonical space and 4D deformation field of a dynamic, non-rigid scene with explicit and discrete voxel-based representations.
 Several recent works extend the above approaches to 3D Gaussian splatting.
@@
 We build on another recent work, 4D Gaussian splatting [53], which uses feature encoding techniques proposed in HexPlanes [6] and K-planes [17], and learns a deformation field instead.
-To the best of our knowledge, all prior deformation-based work learned deformations in a non-metric canonical space without reconstructing metric deformations.
-We propose a method to learn metric deformations from a set of Gaussians with canonical properties, deformed to metric space using a neural-voxel deformation function.
-By training on adjacent timesteps, MD-Splatting allows enforcing physics-informed regularization terms to improve performance beyond photometric consistency.
 2.3 Point Tracking
@@
 Prior work on point tracking often studies tracking 2D points across video frames, where a dominant approach is training models on large-scale synthetic datasets containing ground-truth point trajectories [12, 14, 57, 20] or dense optical flows [48].
-Optical flow [3, 51] or scene flow [56, 2, 58] can also be viewed as single-step point-tracking in 2D and 3D, respectively.
+Optical flow [2, 43] or scene flow [45, 1, 46, 19] can also be viewed as single-step point-tracking in 2D and 3D, respectively.
 Another relevant line of work tightly couples dynamic scene reconstruction and motion estimation of non-rigid objects.
@@
 While our work also leverages 3D Gaussians, in contrary to the explicit modeling in Dynamic 3D Gaussians [29], we learn a deformation function that scales much better with video length, and we focus on deformable objects that are more challenging than the ball-throwing videos used in [29].
-A main motivation for studying point tracking is the potential it can unlock for robotics applications: for example, RoboTAP [55] shows pre-trained point-tracking models improves sample efficiency of visual imitation learning.
+2.4 Tracking for Robotics
+A core motivation for studying point tracking is the potential it can unlock for robotics applications: for example, RoboTAP [44] shows pre-trained point-tracking models improve sample efficiency of visual imitation learning.
+It detects task-relevant keypoints, infers where those points should move to, and computes an action that moves them there.
+Any-point [52] learns to predict keypoint tracks, but conditioned on language inputs.
+Track2Act [3] builds on Any-point by learning a generalizable zero-shot policy, which only needs a few embodiment-specific demonstrations.
+Rigid-body, or 6D, pose tracking and estimation has a rich history in robotics due to it foundational ability to model the world for a robot to manipulate [49, 31, 10, 11, 27, 42, 50, 51, 55].
+In this work, we propose a deformable object analog of 6D pose tracking with the aim of extending successes to deformable object manipulation.
+While existing methods leverage 2D tracking, and learn an additional policy to output robot actions, DeformGS provides a more powerful representation that allows for reasoning directly in 3D, instead of in the 2D image space.
 3 Problem Statement
@@
 Each image sequence is captured over the same time interval $t\in[0,H]$.
-3D Tracking The 3D tracking objective is to recover the trajectory of any point in a dynamic scene by modeling metric deformations.
-Thus, the objective is to find a function $x_{t}=Q(x_{0},t_{0},t)$, where $x_{0}\in\mathbb{R}^{3}$ is the location of a point of interest at a chosen time $t_{0}$, while $x_{t}\in\mathbb{R}^{3}$ is the location of the same point at another chosen time $t$.
+3D Tracking The primary goal is to recover the trajectory of any point in a dynamic scene by modeling the deformation of Gaussians over time.
+Thus, the objective is to find a function $x_{t}=Q(x_{0},t_{0},t)$, where $x_{0}\in\mathbb{R}^{3}$ is the location of a point of interest at a chosen time $t_{0}\in[0,H]$, while $x_{t}\in\mathbb{R}^{3}$ is the location of the same point at another chosen time $t\in[0,H]$.
 The function $Q$ is valid for any point $x_{0}$ and any $t\in[0,H]$, allowing for tracking of any point in space.
-Novel View Synthesis This objective is to recover novel views from arbitrary viewpoints.
+Novel View Synthesis The secondary goal is to achieve accurate scene flow by using photometric consistency as a supervision signal.
+To achieve this, the objective is to recover novel views from arbitrary viewpoints.
 The extrinsics at any viewpoint can be captured by matrix $P$, with $P=K[R|T]$.
 Here $K$ is the intrinsics matrix, $R$ is the rotation matrix of a camera with respect to the world frame, and $T$ is the translation vector with respect to the world frame.
-The aim is to learn a function $V$ such that $I_{P,t}=V(P,t)$, where $I_{P,t}$ is an image rendered from a camera with extrinsics $P$ at time $t$.
+Concretely, the goal is to learn a function $V$ such that $I_{P,t}=V(P,t)$, where $I_{P,t}$ is an image rendered from a camera with extrinsics $P$ at time $t$.
 As with the tracking objective, the time parameter is valid for any $t\in[0,H]$.
 4 Preliminary
-4.1 Neural Radiance Fields (NeRF)
-Neural radiance fields (NeRFs) trace rays through multiple camera views to train an MLP $\Phi_{r}$ to infer a radiance (RGB) and density ($\sigma)$ from a world coordinate $(x)$ and view direction $(\theta,\phi)$ inputs, which are typically concatenated with a positional encoding in the form of sinusoids with varying frequencies.
-$\Phi_{r}(x,\theta,\phi)=(RGB\sigma)$ (1)
 4.1 Gaussian Splatting
-3D Gaussian Splatting [25] deploys a more explicit scene representation than NeRFs, by rendering a large set of Gaussians each defined by their mean position $\mu$ and covariance matrix $\Sigma$.
-Thus, for each $x\in\mathbb{R}^{3}$, its Gaussian $G(x)$ is
+3D Gaussian Splatting [22] deploys an explicit scene representation by rendering a large set of Gaussians each defined by their mean position $\mu$ and covariance matrix $\Sigma$.
+Given $x\in\mathbb{R}^{3}$, its Gaussian is
 $G(x)=e^{-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)},$
-Directly optimizing the covariance matrix $\Sigma$ would lead to infeasible covariance matrices, as they have to be positive semi-definite to have a physical meaning.
+Directly optimizing the covariance matrix $\Sigma$ would lead to infeasible covariance matrices, as they must be positive semi-definite to have a physical meaning.
 Instead, Gaussian Splatting [22] proposes to decompose $\Sigma$ into a rotation $R$ and scale $S$ for each Gaussian:
@@
 $C=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$
-where $c_{i}$ is the color of each point and $\alpha_{i}$ is given by evaluating a 2D Gaussian with covariance multiplied with a learned per-point opacity [67, 25].
+where $c_{i}$ is the color of each Gaussian and $\alpha_{i}$ is given by evaluating a 2D Gaussian with covariance multiplied with a learned per-Gaussian opacity $\sigma$ [56, 22].
+This representation allows for fast rendering of novel views, and aims to reconstruct the geometry of the scene.
 4.2 Deformation Fields for Dynamic Scenes
@@
 [53] applied a similar approach to arrive at Gaussian splatting of dynamic scenes.
-Given the state of a single canonical Gaussian, defined by $P=[\mu,S,R,\sigma,C]$ at time $t$, a deformation function takes the form
+Given the state of a single canonical Gaussian, defined by $P=[\mu,S,R,\sigma,C]$ at time $t$, a deformation function is
 $P^{\prime}=F_{\mathrm{4DGS}}(P,t),$
@@
 During the fine stage we learn the deformation function.
-We propose MD-Splatting (Figure 2), based on 4D-Gaussians [64], to render novel views in dynamic scenes.
-The key differences with 4D-GS are: (1) we propose a simple method to track canonical Gaussians in 3D using a canonical $\rightarrow$ metric deformation function, (2) the output of the deformation function is different, and (3) using the method shown in Figure 3, we enforce regularization losses on the 3D trajectories of Gaussians.
-Figure 2: MD-Splatting maps a set of Gausians with canonical properties to metric space using a deformation function $F$.
+We propose DeformGS (Figure 2), based on 4D-Gaussians [53], to render novel views in dynamic scenes.
+The key differences with 4D-GS are: (1) we propose an intuitive method to track canonical Gaussians in world coordinates using a continuous deformation function, (2) the output of the deformation function is different, e.g., DeformGS infers shadows and does not alter opacity or scale over time, and (3) using the method shown in Figure 3, we enforce physics-inspired regularization losses on the 3D trajectories of Gaussians.
+Figure 2: DeformGS maps a set of Gausians with canonical properties to metric space using a deformation function $F$.
 The deformation function takes in the position of a Gaussian $(x,y,z)$ and a queried timestamp $t$, to infer shadow $s$, rotation $r^{\prime}$ and metric position $x^{\prime}$.
@@
 5 Method
-MD-Splatting achieves high-quality novel view synthesis using a canonical space of Gaussians and a deformation function to deform them to metric space (Section 5.1).
-To incentivize learning physically plausible deformations, MD-Splatting introduces several regularization terms (Section 5.2).
-Figure 3: MD-Splatting uses three adjacent timesteps at every iteration to enforce physics-inspired regularization terms.
-All Gaussians are deformed to the metric space using the deformation function $F$, and rasterized to compute the photometric loss and its gradients.
-The positions of the Gaussians are used to compute the regularization terms based on local rigidity, isometry, and conservation of momentum (Section 5.2).
+DeformGS achieves novel-view synthesis and high-quality 3D tracking using a canonical space of Gaussians and a deformation function to deform them to world space (Section 5.1).
+To incentivize learning physically plausible deformations, DeformGS introduces several regularization terms (Section 5.2).
+Finally, DeformGS learns 3D masks to focus regularization and Gaussian deformation on dynamic parts of the scene (Section 5.3).
 5.1 4D Gaussian Splatting
 Canonical Neural Voxel Encoding.
-As with prior work, MD-Splatting learns a deformation function $F$ from a canonical space.
+As with prior work, DeformGS learns a deformation function $F$ from a canonical space.
 We use a neural-voxel encoding to ensure $F$ has sufficient capacity to capture complex deformations.
-Prior work [17, 64, 16, 36] showed that neural-voxel encodings improve the speed and accuracy of novel view synthesis in dynamic scenes.
-We leverage HexPlanes [6, 64] to increase capacity for metric tracking and novel view synthesis simultaneously.
+Prior work [17, 53, 16, 32] showed that neural-voxel encodings improve the speed and accuracy of novel-view synthesis in dynamic scenes.
+We leverage HexPlanes [6, 53] to increase capacity for simultaneous 3D tracking and novel-view synthesis.
 Figure 2 shows an overview of the canonical neural-voxel encoding.
@@
 Deformation MLP.
-The deformation MLP takes in the voxel encoding and uses the encoding to deform the canonical Gaussians into metric space.
-Figure 2 shows the deformation MLP, which infers position, rotation, and a shadow scalar.
+The deformation MLP takes in the voxel encoding and uses the encoding to deform the canonical Gaussians into world coordinates.
+Figure 2 shows the deformation MLP, which infers position, rotation, and a shadow scalar, given a feature vector from the neural voxel encoding.
 We choose this set of outputs to model rigid-body transformations of each Gaussian and changes in illumination.
 Modeling changes in illumination is critical in the presence of shadows.
-The RGB color of each Gaussian is multiplied by the shadow scalar $s\in[0,1]$, and the shadow scalar is in the range $[0,1]$ by feeding the output of the MLP through a sigmoid activation function.
-Next, we deform the Gaussians, modifying their mean positions $\mu$, rotation $R$, and color $C$, and arrive at a set of Gaussians in the metric space each with state $P$.
-The differentiable rasterizer from Gaussian Splatting [25] then renders the Gaussians to retrieve gradients for regressing both the canonical Gaussian states, and the parameters of the deformation function.
+We multiply the RGB color of each Gaussian by the shadow scalar $s\in[0,1]$, and the shadow scalar is in the range $[0,1]$ by feeding the output of the MLP through a sigmoid activation function.
+Next, we deform the Gaussians, modifying their mean positions $\mu$ and rotation $R$, and arrive at a set of Gaussians in the world space each with state $P$.
+The differentiable rasterizer from Gaussian Splatting [22] then renders the Gaussians to retrieve gradients for regressing both the canonical Gaussian states and the parameters of the deformation function.
 Unlike 4D-Gaussians [53], we propose to not infer opacity or scale using the deformation field.
 Optimizing for opacity and scale over time would allow Gaussians to disappear or appear instead of following the motion, which would make tracking less accurate.
-This design choice reduces the capacity of the deformation function, hence a lower view reconstruction quality as compared to 4D-Gaussians might be expected, see Table 2 for more information.
+This design choice reduces the capacity of the deformation function, hence a lower view reconstruction quality as compared to 4D-Gaussians might be expected.
 5.2 3D Tracking using 4D Gaussians
-Motivation.
-Figure 3 shows the process of tracking Gaussians from the canonical space in metric space.
+Figure 3: DeformGS uses three adjacent timesteps at every iteration to enforce physics-inspired regularization terms.
+All Gaussians are deformed to world space using the deformation function $F$, and rasterized to compute the photometric loss and its gradients.
+The positions of the Gaussians are used to compute the regularization terms based on local isometry and conservation of momentum (Section 5.2).
+Physics-Inspired Losses.
+Figure 3 shows the process of tracking Gaussians from the canonical space in world space.
 By querying the deformation function $F$, we can track the position of a Gaussian along the entire trajectory.
-Without additional supervision, this approach will not necessarily converge to metric deformations.
-Especially when objects include areas with little texture and uniform color, the solution space for all metric deformations is underconstrained by photometric consistency supervision.
-To better learn a metric deformation, we propose regularization terms inspired by physics.
-After empirically evaluating several combinations of regularization terms, we adopt the rigidity and isometry losses proposed in Luiten et al.
-and add a momentum term.
-The first two terms capture a local rigidity loss and local isometry loss, which we compute based on the state of the $k$ nearest neighboring (KNN) Gaussians.
-Local Rigidity Loss We assume the deformable material behaves linearly locally.
-With a large number of Gaussians, dense in 3D space, we expect nearest neighbors to move in similar directions and undergo similar rotations.
-Without this assumption, the trajectories of Gaussians are free to intersect and cross without a penalty.
-Instead, we introduce a local rigidity loss,
-$\mathcal{L}_{i,j,t}^{\text{rigid}}=w_{i,j}\|(\mu_{j,t-1}-\mu_{i,t-1})-R_{i,t-1}R_{i,t}^{-1}(\mu_{j,t}-\mu_{i,t})\|_{2},$
-where $\mu_{i,t}$ is the mean position of Gaussian $i$ at time $t$, $R_{i,t}$ is the rotation matrix of Gaussian $i$ at timestep $t$.
-This loss regularizes the change in position of the $k$ nearest neighbors as seen from Gaussian $i$, by transforming the current offset of the neighbors to the previous Gaussian orientation.
-We define
-$w_{i,j}=\text{exp}\left(-\lambda_{w}\|\mu_{j,0}-\mu_{i,0}\|_{2}^{2}\right),$
-an isotropic Gaussian used to decrease the weight of Gaussians at a greater distance from $\mu_{i,0}$.
-When a sufficiently large $\lambda_{w}$ is applied, and with a small enough $k$ for KNN, this loss fits the local linearity assumption.
-The total rigidity loss at time $t$ is
-$\mathcal{L}^{\text{rigid}}_{t}=\frac{1}{k|\mathcal{P}|}\sum_{i\in\mathcal{P}}\sum_{j\in\text{knn}_{i}}\mathcal{L}_{i,j,t}^{\text{rigid}}.$
-Here $\mathcal{P}$ is the set of all Gaussians.
-Local Isometry Loss Another assumption we incentivize is to keep the relative position of the $k$ nearest neighbors constant w.r.t.
+Without additional supervision, this approach will not necessarily converge to physically plausible deformations.
+Especially when objects include areas with little texture and uniform color, the solution space for all deformations is underconstrained by photometric consistency alone.
+To learn a more grounded deformation function, we propose regularization terms inspired by physics.
+After empirically evaluating several combinations of regularization terms, we adopt the isometry loss proposed in [29] and add a conservation of momentum term.
+The first term captures a local isometry loss, which we compute based on the state of the $k$ nearest neighboring (KNN) Gaussians.
+Local Isometry Loss We incentivize the Gaussians to keep the relative position of the $k$ nearest neighbors constant w.r.t.
 $t=0$.
-With sufficient deformation, this assumption will be broken at a larger scale, but at a local scale, this regularization avoids drift from the ground truth trajectory.
+With sufficient deformation, this assumption will be broken at a larger scale, but at a local scale, this regularization avoids drift from the ground-truth trajectory.
 The isometry loss is
 $\mathcal{L}_{t}^{\text{iso}}=\frac{1}{k|\mathcal{P}|}\!\sum_{i\in\mathcal{P}}\!\sum_{j\in\text{knn}_{i}}\!\!w_{i,j}\!\left|\left\|\mu_{j,0}{-}\mu_{i,0}\right\|_{2}-\left\|\mu_{j,t}{-}\mu_{i,t}\right\|_{2}\right|.$
-Conservation of Momentum The final term we add is to incentivize the conservation of momentum.
+with
+$w_{i,j}=\text{exp}\left(-\lambda_{w}\|\mu_{j,0}-\mu_{i,0}\|_{2}^{2}\right),$
+Here $\mathcal{P}$ is the set of all Gaussians.
+Conservation of Momentum We add a term to incentivize conservation of momentum.
 Newton’s first law states objects without external forces applied, given some mass $m$ and velocity vector $\mathbf{v}$, maintain their momentum $m\cdot\mathbf{v}$.
@@
 $\mathcal{L}^{\text{momentum}}_{i,t}=\|\mu_{i,t+1}+\mu_{i,t-1}-2\mu_{i,t}\|_{1}.$
-This term incentivizes a constant velocity vector and has the effect of imposing a low-pass filter on the 3D trajectories.
+This term incentivizes a constant-velocity vector and has the effect of imposing a low-pass filter on the 3D trajectories.
 It smooths out trajectories with many sudden changes of direction and magnitude (momentum).
-Table 1 shows the impact of the proposed regularization terms on tracking errors.
+5.3 Learning 3D Masks
+Learning accurate 3D tracking in scenes with a mix of static and dynamic objects and rich textures poses significant challenges, mainly: (1) imposing physics-inspired regularization terms on all Gaussians may cause issues when dynamic and static objects interact, and (2) modeling millions of dynamic Gaussians can become a significant computational burden.
+To address this, DeformGS takes noisy masks of dynamic scene components such as cloth, and learns what Gaussians are dynamic. More formally, we render a mask $M$ by
+$M=\sum_{i\in N}m_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$
+where $m_{i}$ is a per-Gaussian property, with $m_{i}\in[0,1]$.
+We then add a regularization term to the loss function s.t.
+$m_{i}$ is regressed to best reconstruct M.
+Finally, DeformGS uses $m_{i}$ to select a subset of Gaussians to be dynamic, and applies regularization terms only to those Gaussians.
 6 Experiments
-We test the effectiveness of MD-Splatting on synthetic datasets of deformable objects.
-Section 6.1 provides details on our experiment setup, evaluation metrics, and baseline methods.
-Section 6.2 reports evaluation results from the compared methods and result analysis.
-Section 6.3 supplies additional ablation results on the training efficiency of MD-Splatting.
-6.1 Experiment Setup
-Metric Method Scene 1 Scene 2 Scene 3 Scene 4 Scene 5 Scene 6 Mean
-3D MTE [mm] $\downarrow$ RAFT [51] ${}^{a}$ 89.687 151.363 295.66 135.468 125.554 67.861 144.266
-RAFT Oracle [51]${}^{a}$${}^{b}$ 27.328 34.629 104.842 24.976 32.331 36.113 43.370
-OmniMotion Oracle [60] ${}^{a}$${}^{b}$ 16.188 23.51 92.536 11.064 57.603 19.624 36.754
-DeVRF [30] 147.293 78.487 449.74 123.745 77.736 133.391 168.399
-DynaGS [33] 17.189 54.409 61.148 19.870 24.081 26.757 33.909
-4D-GS [64] 7.788 93.409 118.692 16.422 7.033 4.600 41.324
-MD-Splatting-NV 5.173 63.894 81.917 9.449 5.679 3.385 28.250
-MD-Splatting-TK 2.881 46.257 88.169 9.686 4.741 3.175 25.818
-3D $\delta_{\mathrm{avg}}~{}\uparrow$ RAFT [51] ${}^{a}$ 0.442 0.229 0.128 0.295 0.395 0.434 0.321
-RAFT Oracle [51] ${}^{a}$${}^{b}$ 0.623 0.562 0.247 0.651 0.573 0.509 0.528
-OmniMotion Oracle [60] ${}^{a}$${}^{b}$ 0.763 0.612 0.301 0.763 0.448 0.680 0.595
-DeVRF [30] 0.344 0.368 0.094 0.326 0.471 0.398 0.334
-DynaGS [33] 0.774 0.401 0.317 0.697 0.649 0.607 0.574
-4D-GS [64] 0.787 0.337 0.257 0.700 0.817 0.883 0.630
-MD-Splatting-NV 0.851 0.428 0.346 0.749 0.870 0.919 0.694
-MD-Splatting-TK 0.912 0.511 0.338 0.780 0.903 0.939 0.731
-3D Survival $\uparrow$ RAFT [51] ${}^{a}$ 0.935 0.891 0.738 0.919 0.807 0.891 0.864
-RAFT Oracle [51] ${}^{a}$${}^{b}$ 1.000 1.000 1.000 1.000 0.999 0.974 0.996
-OmniMotion Oracle [60] ${}^{a}$${}^{b}$ 1.000 0.993 1.000 0.991 0.870 1.000 0.976
-DeVRF [30] 1.000 0.972 0.571 0.952 0.860 0.862 0.870
-DynaGS [33] 1.000 1.000 0.958 1.000 1.000 1.000 0.993
-4D-GS [64] 1.000 0.978 0.967 1.000 0.999 1.000 0.991
-MD-Splatting-NV 0.999 1.000 0.987 1.000 1.000 1.000 0.998
-MD-Splatting-TK 1.000 1.000 1.000 1.000 1.000 1.000 1.000
-${}^{a}$ Method had access to ground truth depth. ${}^{b}$ Method had access to ground truth trajectories to pick the best camera view.
-Table 1: 3D tracking results on the deformable cloth dataset (Figure 1).
-For each metric, the methods above the solid line had access to privileged information, see ${}^{a}$${}^{b}$ and Section 6.1 for more details.
-The results suggest that MD-Splatting outperforms the baselines in all averaged metrics.
-DynaGS [33] and OmniMotion [60] are both competitive but accumulate large tracking errors in scenes 1, 5, and 6.
-The results also show the regularization parameters improve the tracking performance compared to 4D-GS [64].
+We evaluate DeformGS on synthetic and real-world datasets of scenes with highly deformable objects.
+Section 6.1 provides details on the simulation experiment setup, evaluation metrics, and baseline methods.
+Section 6.2 reports evaluation results from the compared methods and provides analysis.
+Section 6.3 lists the real-world evaluation setup, and finally in Section 6.4 we provide a qualitative evaluation of the performance of DeformGS in the real-world.
+6.1 Simulation Experiment Setup
 Dataset Preparation We use Blender to model dynamic cloth sequences and render photo-realistic images.
@@
 We obtain ground-truth trajectories by tracking the mesh vertices of deformable objects in Blender.
-Baselines.
-We compare to baseline methods: (1) 2D tracking approaches by projecting their inferences to 3D points using Blender’s ground truth depth and the camera’s ground truth projective geometry.
+Every scene contains a single deformable object and a rendered background.
+Oracle Baselines We compare DeformGS to 2D tracking oracle models which have access to ground truth depth and trajectory information.
 While these methods were not designed for 3D tracking, they are well-known for their impressive 2D tracking performance.
 Their numbers aid in putting the tracking performance of the other baselines into context.
-We run RAFT [51] on all views and report the mean results as the RAFT model.
-We also provide two oracle methods which have access to privileged information: the ground truth trajectories.
+We run RAFT [43] on all views, project tracking to 3D using ground truth depth, and report the mean results as the RAFT model.
+We provide two additional oracle methods which have access to the ground-truth trajectories as well.
 RAFT Oracle first evaluates on all views, to then output only the trajectories from the view with the lowest median trajectory error.
-We also report OmniMotion Oracle, which runs OmniMotion [60] on the frame with the lowest MTE for RAFT.
+We also report OmniMotion Oracle, which runs OmniMotion [48] on the viewpoint with the lowest MTE for RAFT.
 Training OmniMotion takes roughly 12–13 hours on an Nvidia RTX 4090 GPU, making inference on all 100 views impractical.
-The numbers from RAFT Oracle and OmniMotion Oracle are not an apples-to-apples comparison with the other baselines, as to obtain their result they had to access privileged ground-truth trajectories.
-(2) DeVRF [30] represents a dynamic, non-rigid scene with a 3D canonical voxel space and a 4D deformation voxel field.
-It proposes a 4D deformation cycle that models both the backward deformation from the deformed space to the canonical space and the forward deformation from the canonical space to the deformed space.
-Therefore, the point trajectories can be obtained by querying the forward deformation of canonical points at $t=0$ to every timestep’s deformed space through the 4D forward deformation space.
-While DeVRF was not designed for 3D tracking directly, it provides useful insight into how similar methods based on NeRF might perform in highly deformable scenes.
-(3) Dynamic 3D Gaussians (DynaGS[33], which also builds upon 3D Gaussian splatting for dynamic novel view synthesis, except it explicitly models the positions and rotations of each Gaussian at each time-step.
-This results in straightforward tracking of any point via finding the trajectory of the learned Gaussian with the highest influence on the given point.
-Although the original paper assumes a known point cloud at the first frame, we instead use a randomly sampled point cloud for a fair comparison, with DynaGS and MD-Splatting both not using depth information.
-We also ran with a point cloud initialized using ground-truth depth from Blender but saw very little to no improvement.
+The numbers from RAFT Oracle and OmniMotion Oracle are not an apples-to-apples comparison with the other methods, as to obtain their result they had to access privileged ground-truth trajectories.
+Gaussian Splatting Baselines (1) Dynamic 3D Gaussians (DynaGS[29], which also builds on 3D Gaussian splatting for dynamic novel-view synthesis, except it explicitly models the positions and rotations of each Gaussian at each time-step.
+This results in straightforward tracking of any point via finding the trajectory of the learned Gaussian closest to a queried point.
+Although the original paper assumes a known point cloud at the first frame, we instead use a randomly sampled point cloud for a fair comparison, with DynaGS and DeformGS both not using depth information.
 (2) Finally, we compare to tracking using 4D-Gaussians [53] (4D-GS).
-We add our approach for tracking canonical Gaussian, as shown in Figure 3, to extract 3D trajectories from a learning view synthesis model.
-Comparing to 4D-GS serves to show the impact of the changes made in the model architecture and the regularization terms, to arrive at MD-Splatting.
-Metrics Method Scene1 Scene2 Scene3 Scene 4 Scene5 Scene 6 Mean
-[…]
-DynaGS [33] 35.51 32.23 31.36 30.54 29.84 30.48 31.66
-4D-GS [64] 40.76 40.14 44.38 42.05 33.70 34.83 39.31
-MD-Splatting-NV 40.67 39.62 43.27 42.61 33.46 34.96 39.10
-MD-Splatting-TK 38.76 38.16 40.34 40.61 32.47 32.39 37.12
-[…]
-DynaGS [33] 0.96 0.93 0.95 0.96 0.92 0.95 0.94
-4D-GS [64] 0.99 0.97 0.98 0.99 0.96 0.98 0.98
-MD-Splatting-NV 0.99 0.97 0.98 1.00 0.96 0.98 0.98
-MD-Splatting-TK 0.99 0.97 0.98 0.99 0.95 0.97 0.98
-[…]
-DynaGS [33] 0.03 0.11 0.16 0.07 0.10 0.06 0.09
-4D-GS [64] 0.01 0.07 0.07 0.02 0.07 0.02 0.04
-MD-Splatting-NV 0.01 0.07 0.07 0.02 0.08 0.02 0.05
-MD-Splatting-TK 0.02 0.08 0.08 0.03 0.08 0.04 0.05
-Table 2: Novel view synthesis results on the deformable cloth dataset (Figure 1).
-The results suggest MD-Splatting significantly outperforms DynaGS [33] and DeVRF [30], while 4D-GS [64] performs slightly better.
-This is expected as a result of the additional regularization terms and deformation function design (Section 5).
-MD-Splatting only takes a minor decrease in view reconstruction performance, in order to significantly improve tracking.
-The results also suggest that MD-Splatting-TK performs worse in tracking as compared to MD-Splatting-NV, due to the stronger isometry regularization (Section 6.4).
+We add the approach for 3D tracking of canonical Gaussian, as shown in Figure 3, to extract 3D trajectories from a learning view synthesis model.
+Comparing to 4D-GS serves to show the impact of the changes made in the model architecture, the regularization terms, and using learning per-Gaussian masks to arrive at DeformGS.
 […]
 […]
-We evaluate 3D tracking and novel-view synthesis results.
-For 3D tracking we evaluate each compared method on 1,000 randomly sampled points on each cloth.
-For novel-view synthesis, we evaluate using 200 unseen test-time cameras, and average the results across all cameras and all 40 time steps.
-Figure 4: Qualitative novel view synthesis results from all baselines and MD-Splatting.
-Both 4D-GS [64] and MD-Splatting achieve the best qualitative results that are close to ground-truth renderings (especially on Scene 6, where the floral patterns on the cloth are reconstructed), albeit still fail to capture the fine-grained texture and wrinkles on Scene 2 and Scene3 cloth.
-6.2 Main Results
-3D Point Tracking. Following prior work [71, 33], we report median trajectory error (MTE), position accuracy ($\delta$), and the survival rate with a threshold of 0.5 [m] [33].
+We evaluate each compared method on 1,000 randomly sampled points on each cloth.
+6.2 Simulation Results
+Metric Method Scene 1 Scene 2 Scene 3 Scene 4 Scene 5 Scene 6 Mean
+3D MTE [mm] $\downarrow$ RAFT [43] a 67.264 89.944 220.125 177.909 84.593 23.422 110.543
+RAFT Oracle [43]ab 3.381 26.956 58.971 12.481 3.930 3.192 18.152
+OmniMotion Oracle [48] ab 0.535 14.513 39.958 4.556 2.487 2.011 10.677
+DynaGS [29] 24.233 81.119 464.64 54.074 34.985 36.101 115.859
+4D-GS [53] 4.645 95.032 223.839 27.441 14.091 11.619 62.778
+DeformGS 3.373 45.33 88.369 14.022 7.257 8.173 27.754
+3D $\delta_{\mathrm{avg}}~{}\uparrow$ RAFT [43] a 0.553 0.379 0.222 0.533 0.586 0.619 0.482
+RAFT Oracle [43]ab 0.926 0.577 0.411 0.703 0.833 0.808 0.710
+OmniMotion Oracle [48] ab 0.987 0.693 0.549 0.864 0.885 0.849 0.805
+DynaGS [29] 0.624 0.31 0.042 0.435 0.533 0.527 0.4118
+4D-GS [53] 0.902 0.339 0.164 0.583 0.697 0.715 0.5667
+DeformGS 0.929 0.522 0.322 0.71 0.856 0.787 0.688
+3D Survival $\uparrow$ RAFT [43] a 0.945 0.792 0.779 0.822 0.792 0.854 0.831
+RAFT Oracle [43]ab 0.986 0.833 0.957 0.872 0.929 0.903 0.913
+OmniMotion Oracle [48] ab 1 0.963 1 0.985 0.963 0.933 0.977
+DynaGS [29] 0.999 0.99 0.483 0.988 0.992 0.992 0.907
+4D-GS [53] 1 0.967 0.834 1 0.999 1 0.967
+DeformGS 1 1 1 1 1 1 1
+a Method had access to ground truth depth. b Method had access to ground truth trajectories to pick the best camera view.
+Table 1: 3D tracking results on the deformable cloth dataset (Figure 1).
+For each metric, the methods above the solid line had access to privileged information, see ab and Section 6.1 for more details.
+The results suggest that DeformGS outperforms the baselines in all averaged metrics, and is competitive with the oracle models.
+The results also suggest our novel deformation function architecture, learning per-Gaussian masks, and physics-inspired regularization losses improve the tracking performance compared to 4D-GS [53].
+We do not consider the oracle methods to be fair baselines and therefore do not bold their results.
+3D Point Tracking Following prior work [57, 29], we report median trajectory error (MTE), position accuracy ($\delta$), and the survival rate with a threshold of 0.5 [m] [29].
 The results are summarized in Table 1.
-We make the following observations: (1) MD-Splatting outperforms the baselines in all averaged metrics, with a 23.9% better average MTE compared to the baseline methods.
-Dynamic 3D Gaussians [33], OmniMotion [60], and RAFT [51] are competitive but accumulate a significantly larger tracking error in scenes 1, 4, 5, and 6.
-For example, in scene 1, the baselines have a MTE of at least 170% higher as compared to MD-Splatting.
-Excluding our implementation of 4D-GS [64] for tracking, the clostest baseline is the OmniMotion [60] oracle model with a MTE 462% higher as compared to MD-Splatting.
-(2) MD-Splatting-TK performs better tracking compared to MD-Splatting NV, except for scene 3.
-The results suggest that setting $\lambda^{\text{iso}}=\text{1.0}$ leads to overall better tracking performance.
-Section 6.4 contains a more in-depth analysis on this parameter.
-(3) MD-Splatting achieves the best performance in environments with textured cloth: scenes 1, 5, and 6.
-The texture makes photometric consistency a stronger supervision signal, leading to more accurate 3D tracking.
-(4) The regularization parameters and change in deformation function architecture seem to improve the tracking performance compared to our implementation of 4D-GS [64].
-MD-Splatting scores better metrics in all scenes.
-(5) The RAFT [51] and OmniMotion [60] results are highly dependent on the camera view angles, and we report the camera view with the best error values after evaluating all the test-time views; despite this given advantage, these methods still fail to achieve consistent and competitive tracking performance.
-We hypothesize that part of the tracking error from Dynamic 3D Gaussians [33] is a result of converging Gaussian states to local optima.
-Especially with larger deformations and shadows, we imagine Gaussians could converge to local optima with relatively large tracking errors.
-The other source of error we hypothesize plays a role is the various regularization terms and their hyperparameters, which are predominantly focused on rigid motion.
-(a) DeVRF [30]
+We make the following observations: (1) DeformGS outperforms baselines RAFT, DynaGS, and 4D-GS, by achieving a MTE of 55.8% - 76.0% lower compared to the baselines.
+(2) The discrepancy between the RAFT oracle model and its averaged result demonstrates the difficulty arising from frequent self-occlusions.
+This also points to future research avenues for additional supervision through optic flow and 2D tracking algorithms such as RAFT.
+(3) The oracle models perform very well, this is in part thanks to the falling and short-horizon nature of these sequences, limiting self-occlusions.
+In the real-world we expect much larger errors due to noisy depth and more challenging occlusions in long-horizon tasks.
+It would also be unclear what viewpoint to choose without access to ground truth trajectories.
+(4) Scenes with less texture such as scene 3 perform significantly worse than scenes with strong texture.
+Qualitative Results Figure 4 shows ground truth and inferred trajectories in scene 5.
+The results show that especially DynaGS and 4D-GS introduce large errors as the cloth drapes down.
+RAFT improves over DynaGS and 4D-GS but requires accurate depth estimation.
+(a) RAFT [43]
 (b) DynaGS [29]
 (c) 4D-GS [53]
-(d) MD-Splatting
-Figure 5: Results on Scene 6: randomly sampled ground-truth trajectories in green, inferred trajectories in red, and the error in red.
-The differences between 4D-GS [64] are marked with green squares.
-Compared to the baseline methods, MD-Splatting results in fewer errors in 3D tracking.
-Novel View Synthesis.
-We evaluate visual reconstruction quality.
-In Table 2, we report results on the three standard metrics commonly used in prior work: PSNR [61], SSIM [62], and LPIPS [70].
-MD-Splatting-NV performs comparably to 4D Gaussians (4D-GS) on SSIM while obtaining marginally lower results on PSNR and LPIPS.
-DynaGS performs worse on view synthesis by a considerable margin.
-DeVRF obtains the lowest performance, and it does worse on the Scene 2 and 3 as compared to the other scenes—this suggests the textureless cloths impose additional challenges that the model cannot capture.
-MD-Splatting-TK yields a lower quality view reconstruction as compared to MD-Splatting-NV, as a result of the stronger isometry regularization (Section 6.4).
-We jointly analyze results from 3D tracking 1 and novel view sythesis 2.
-We observe that 4D-GS under-performs DynaGS at 3D tracking in median trajectory error, but outperforms DynaGS at novel view synthesis.
-MD-Splatting is designed to achieve the best of both worlds.
-While close to 4D-GS in nature, MD-Splatting can track deformations better than DynaGS while keeping the high-quality novel-view synthesis capability.
-Qualitative Results: Tracked Trajectories Figure 5 shows a quantitative comparison for tracking performance in Scene 6, overlaid on each method’ renders.
-The randomly sampled trajectories show examples of trajectories where DeVRF [30] and Dynamic 3D Gaussians [33] diverge from the ground truth.
-For DeVRF [30], we hypothesize a lack of high-quality view reconstruction is the cause of poor tracking, whereas Dyamic 3D Gaussians [33] has high-quality view reconstruction but has a few trajectories diverge from the ground truth.
-The difference between 4D-GS and MD-Splatting-NV is more subtle, and can be observed in the areas marked with green boxes.
-The green boxes highlight areas where 4D-GS [64] struggles to maintain tracking, whereas MD-Splatting-NV achieves better tracking.
-Qualitative Results: Novel View Synthesis.
-Figure 4 shows a qualitative comparison for scenes 2, 3, and 6.
-Both 4D-GS and MD-Splatting-NV achieve the best qualitative results that are close to ground-truth (notably on Scene 6, where the floral patterns on the cloth are reconstructed), albeit still fail to capture the fine-grained texture and wrinkles on Scene 2 and 3 cloth.
-DynaGS also achieves high-quality rendering overall but display a blurry visual effect around the deformed (i.e., wrinkled) areas.
-We hypothesize that regressing Gaussian positions and orientations is more unstable, and prone to converge to local optima.
-DeVRF nearly completely fails in Scene 3 due to the lack of textures: the model overfits to one small patch on the cloth and ignores the rest of the scene; it also shows difficulty in learning the large deformations on edges, i.e.
-only capturing colors around the center of the cloth.
-6.3 Training Efficiency Ablation
-We compare the training efficiency between MD-Splatting-NV and Dynamic 3D Gaussians [33] when decreasing the number of time steps provided from each camera view.
-We evaluate on Scene 5 and report PSNR on the subsampled frames, training time, and median trajectory error in Table 3.
-We observe that for MD-Splatting, training time is less affected by the number of frames from each camera view, as compared to Dynamic 3D Gaussians [33].
-This is a result of its approach to train one time step at a time.
-The results also suggest that MD-Splatting has more robustness to larger deformations between frames, leading to a better PSNR and median trajectory error.
-Metric Method 40 Steps 20 Steps 14 Steps 10 Steps
+(d) DeformGS
+Figure 4: Results on Scene 5: randomly sampled ground-truth trajectories in green, inferred trajectories in red, and the error of corresponding points in red lines.
+Compared to the baseline methods, DeformGS results in fewer errors in 3D tracking.
+6.3 Real-World Experiment Setup
+Robo360 Data The Robo360 dataset [26] is a 3D omnispective multi-material robotic manipulation dataset.
+It covers many different scenario’s, including manipulation by robot manipulators and humans captured by 86 calibrated cameras.
+These properties make it an ideal dataset to evaluate the effectiveness of DeformGS in the real world.
+We select two scenes: (1) a human folding a larger duvet and (2) a human folding a smaller cloth.
+In the cloth folding scene, we exclude viewpoints where the entire person’s body is visible to eliminate unnecessary complexity.
+We also subsample the data to demonstrate DeformGS performance with fewer views. The duvet folding scene contains 17 training views and the cloth folding scene contains 20 training views.
+6.4 Real-World Experiment Results
+Real2Sim for Digital Twins Figure 5 shows the 3D tracking overlaid on rendered images, as well as the Gaussian points at each time step.
+The results suggest that DeformGS is able to successfully infer smooth and meaningful trajectories in the real world.
+While no ground truth is available, the trajectories appear to follow their geometry closely except for a few floating Gaussians.
+Hyperparameter tuning of the regularization functions, as well as discarding Gaussians with a low opacity, might help resolve this.
+The point cloud included in this Figure can be used to create a digital twin after recording the sequence.
+The digital twin of the duvet, and the entire environment, can then be used to create more dense supervision for imitation learning approaches.
+Task-Relevant Keypoint Tracking Robotic manipulators can benefit from tracking task-relevant keypoints, such as the corner of a cloth or the edge of a jacket.
+Figure 6 shows a comparison between 4D-GS [53] and DeformGS in 3D point tracking, evaluated on both duvet and cloth scenes.
+The results suggest DeformGS leads to more smooth and overall useful trajectories.
+The trajectories from 4D-GS intertwine into more messy trajectories, and appear less physically plausible.
+This would hinder the adoption of 3D tracking into robot applications.
+Figure 5: A person manipulating a duvet in the Robo360 [26] dataset, reconstructed using DeformGS.
+The top row shows the 4D Gaussians as point clouds, where the color represents dense correspondences.
+The bottom row shows rendered views overlaid with 3D trajectories projected to image space.
+Figure 6: Real-world results comparing our proposed DeformGS against 4D-GS [53].
+The 3D trajectories inferred by DeformGS appear more smooth and accurate, whereas 4D-GS displays more cluttered trajectories.
+7 Conclusions
+In this work, we address the challenging problem of 3D point-tracking in dynamic scenes with deformable objects.
+We introduced DeformGS, the first approach that learns continuous deformations for 3D tracking of deformable scenes.
+We empirically demonstrate that DeformGS outperforms baseline methods and achieves both high-quality dynamic scene reconstruction and high-accuracy 3D tracking on highly deformed cloth objects with occlusions and shadows, both in simulation and the real world.
+We also contribute a dataset of six synthetic scenes to facilitate future research.
+Limitations and Future Work DeformGS, similar to prior work on dynamic novel view reconstruction, requires a setup of multiple synchronized and calibrated cameras, which may require a significant engineering effort in real-world scenarios.
+Additionally, significant innovation will be required to achieve the demonstrated results in real-time, as will be beneficial for scalable robot applications.
+While DeformGS improves upon prior methods, we do observe Gaussians wandering off in some cases.
+We also notice the algorithm is relatively sensitive to the regularization hyper parameters ($\lambda^{\text{momentum}}$ and $\lambda^{\text{iso}}$), this might be resolved in the future by adding supervision from state-of-the-art point-tracking algorithms.
+These limitations point to promising directions for future research.
+Acknowledgements This work was supported by the Center for Machine Learning and Health (CMLH) at CMU, and the Pittsburgh Super Computing Center (PSC).
+We thank David Held for the productive discussions.
+{credits}
 […]
-MD-Splatting 33.47 33.00 32.80 32.91
-Training DynaGS [33] 94.96 81.71 71.43 61.30
-Time [mins] $\downarrow$ MD-Splatting 35.68 21.12 20.21 23.76
-MTE [mm]$\downarrow$ DynaGS [33] 28.48 43.04 65.74 110.35
-MD-Splatting 5.789 6.725 8.876 7.876
-Table 3: Ablation on the number of time steps provided for each camera view. Fewer time steps means a lower time resolution.
-[…]
-6.4 Isometry Loss Ablation
-[…]
-7 Conclusions
-In this work, we address the challenging problem of simultaneous 3D point-tracking and novel view synthesis in dynamic scenes with highly deformable objects.
-We introduced MD-Splatting, the first approach that learns continuous metric deformations from a canonical space of Gaussians.
-We empirically demonstrate that MD-Splatting outperforms baseline methods and achieves both high-quality dynamic scene reconstruction and high-accuracy 3D tracking on highly deformed cloth objects with occlusions and shadows.
-We also contribute a dataset of six synthetic scenes to facilitate future research.
-Limitations and future work.
-MD-Splatting, similar to prior work on dynamic novel view reconstruction, requires a setup of multiple synchronized and calibrated cameras, which may require a significant engineering effort in real-world scenarios.
-This work also focused on scenes with only a cloth as the visible object, more scenes, simulation or real, with soft objects in a larger environment would be helpful.
-These limitations point to promising directions for future research.
+The authors have no competing interests to declare that are relevant to the content of this article.